The probability of being choosen for a dance committee is 3/40. There are 360 students in your class. How many people will be on the dance committee?
A. 9
B. 27
C. 40
D. 120

Answer by checkley71(8403)

3/40=X/360
40X=3*360
40X=1080
X=1080/40
X=27 NUMBER OF PEOPLE ON THE DANCE COMMITTEE.